# Do Mice Grok? Glimpses of Hidden Progress During Overtraining in Sensory Cortex

### v1

@@ 4.2 Rich Learning During Overtraining can Explain Overtraining Reversal
A Mathematical Model. We show one sense in which having better features leads to easier reversal. Our argument will relies solely on a notion of feature learning that equates to higher alignment between the task function and model weights. This increased alignment between a one model overtrained longer than the other is analogous to the difference in alignment between a randomly initialized model and a pretrained model, and this is the setting we will reason about. Specifically, we argue that a model pretrained on a task $y(x)$ will learn its reversal $-y(x)$ faster than a randomly initialized model because it has features more aligned to the task, and this increased alignment is exactly the key difference between a model trained to saturation in train accuracy, and one overtrained long long past this saturation point.
Let $\phi(x)$ denote the hidden representations of a neural network trained to learn a function $y(x)$. Consider training of the readout weights $w$ on the reversal task $-y(x)$ with the either pretrained features $\phi(x)$ or with the initial random features $\phi_{0}(x)$. It is known that one effect of feature learning is an increased alignment between the final feature kernel $K(x,x^{\prime})=\phi(x)\cdot\phi(x^{\prime})$ and the target function $y(x)$ (Bordelon and Pehlevan, 2022). Let the original (random) features be given by feature kernel $K_{0}$ (for instance, the initial neural tangent kernel of a neural network). Then letting $y$ represent the labels in the dataset and assuming $(\bm{K},\bm{K}_{0})$ represent the pretrained and original kernel matrices (respectively) evaluated on the training dataset for the original task $y(x)$, we can write
$y^{\top}\exp(-2\bm{K}t)y\leq y^{\top}\exp\left(-2\bm{K}_{0}t\right)\ y\quad\forall t\geq 0.$ (2)
This condition is a common measure of task-model alignment in machine learning Kornblith et al. (2019); Cortes et al. (2012), and can be achieved in many ways. If in both cases, $y$ is an eigenvector of $K$ and $K_{0}$, then this condition is equivalent to the inequality $y^{\top}\bm{K}y\geq y^{\top}\bm{K}_{0}y$. In general, improved alignment of top eigenvectors of $K$ with $y$ can also help achieve this inequality. Feature learning has been empirically observed to increase alignment through both of these effects in multilayer neural networks, see (Canatar and Pehlevan, 2022; Bordelon and Pehlevan, 2022). This formalism merely captures the intuitive notion that “feature learning” in a neural network can be seen as adapting weights to task-relevant directions to allow for generalization gains (compared to lazy training, where a neural network behaves like a kernel method). Then under gradient descent, training on the reversed task $-y$ with MSE loss, the loss evolves as:
$\mathcal{L}_{\text{pretrained}}=|\left(-y\right)-\bm{f}(t)|^{2}=y^{\top}\exp\left(-2\bm{K}t\right)y\leq y^{\top}\exp\left(-2\bm{K}_{0}t\right)y=\mathcal{L}_{\text{random}}$ (3)
which illustrates that using the pretrained features to learn the reversal task will improved the convergence rate compared to the original features. The intuition is that if features have sufficiently high alignment to the task structure, it suffices to just tune the readouts on frozen features to learn task reversal. This is in contrast to having to learn or modify both features and readouts from scratch in a less aligned (for instance, random) setting. We can see such an advantage concretely by revisiting our synthetic model of piriform cortex. We train our MLP to convergence on the original binary classification task, then flip the task labels $y(x)\mapsto-y(x)$. We then retrain only the readouts of the pretrained model, and a randomly initialized network and find that the model pretrained on the original task $y(x)$ learns the inverted task to low test error faster than the randomly initialized model (Figure 7). We find a random model cannot learn the task to low test error if only its readouts are trained. In the reversal effect as studied by experimental psychologists, this neural explanation posits that an animal overtrained on a task will have better task-model alignment and a richer feature set than an animal whose learning was stopped after behavioral plateau, and hence will adapt to reversal faster because it can repurpose existing features.
Figure 7: The model exhibits overtraining reversal: because the pretrained model learns task-relevant features during overtraining, it can re-learn the modified task quickly by just modifying its readout values. In contrast, a randomly initialized network must learn task features from scratch.
@@ A.1 Methodological and implementation details
The raw data consists of around 70 sessions, each representing a recording of  20 random neurons on a given mouse-day. The data is spikes over time for each neuron on each trial, of which there are 200-300 in a session. After odor onset, mice are given 2.5s to lick in response. We count the number of spikes in the final 500ms part of this window and decode on this number, which gives us a $[N,T]$ neurons by trials spike count matrix to decode on for each mouse-day. We drop neurons with low (less than 4) number of spikes across all trials classifying them as unresponsive to the odors, we z-score this matrix and use 10-fold cross validation to get decoding accuracy, over 20 iterations. Since session to session data is very noisy, we plot smoothed statistics using a sliding window and plot standard error of all statistics. It should be noted the last few days of mouse $T$ overtrainining are dropped because of notable inconsistencies in neural data such as the decoding accuracy plummeting to far below chance while behavior remained high, suggesting an unknown and idiosyncratic cause for unreliable decoding accuracy in this mouse. [^2] Berners-Lee et al. [2023] include these unreliable decoding trials in their analysis. Note that the activity matrix was z-scored across trials, so every neuron contributed equally in our analysis.
@@ A.2.2 Important Experimental Details for General Sensory Discrimination Tasks
A.3 Mathematical Details on Fisher Discriminant, LDA, and Margin
@@ A.4 Ablations
A.5 Continued Introduction

### v2

@@ 4.2 Rich Learning During Overtraining can Explain Overtraining Reversal
A Mathematical Model. We illustrate this with the simplest possible network model that exhibits this effect, a two layer linear network (Saxe et al., 2013). Reversal learning proceeds in two steps. First, we train a two layer linear network with $N$ hidden neurons $f(\bm{x})=\frac{1}{N\gamma_{0}}\bm{w}^{2}\cdot\bm{h}(\bm{x})$, where $\bm{h}(\bm{x})=\bm{W}^{1}\bm{x}$ is optimized for $T$ steps on task $y(x)$. After this, a new readout vector $\bm{v}$ is introduced to give a new output $f_{\text{rev}}(\bm{x})=\frac{1}{N\gamma_{0}}\bm{v}\cdot\bm{h}(\bm{x})$ which is trained on the reversed task $-y(\bm{x})$. The matrix $\bm{W}^{1}$ can be updated in the second phase of learning as well. At large width $N$, we can invoke mean field theory ideas (Bordelon and Pehlevan, 2022, see Appendix F.1.1), allowing us to describe the hidden kernel $K(\bm{x},\bm{x}^{\prime})=\frac{1}{N}\bm{h}(\bm{x})\cdot\bm{h}(\bm{x}^{\prime})$ dynamics in both phases of training. For whitened data and squared error loss, it suffices to compute the projection of the kernel $\bm{K}$, $K_{y}=\bm{y}^{\top}\bm{K}\bm{y}$, and the projection of the predictions $f_{y}=\bm{y}^{\top}\bm{f}$ along the direction of the labels $\bm{y}$, which gives:
$K_{y}(t)=\sqrt{1+\gamma_{0}^{2}f_{y}(t)^{2}}\ ,\ \partial_{t}f_{y}(t)=2\sqrt{1+\gamma_{0}^{2}f_{y}(t)^{2}}\ (y-f_{y}(t)).$ (2)
We show that the kernel evolves only in this $\bm{y}\bm{y}^{\top}$ direction in Figure 7 (a), where points separate by their target category. These dynamics are run from time $t=0$ to time $T$, which gives a terminal value of the kernel alignment $K_{y}(T)$. In the second phase of learning ($t>T$) where the new output $f_{\text{rev}}$ is fit, the predictions on the reversal task in the $\bm{y}$ direction evolve as
$\partial_{t}f_{\text{rev},y}(t)=\sqrt{(K_{y}(T)+1)^{2}+4\gamma_{0}^{2}f_{\text{rev},y}(t)^{2}}\ (-y-f_{\text{rev},y}(t)).$ (3)
We note that larger kernel-target overlap $K_{y}(T)$ leads to faster training in this second phase $t>T$, as rate of learning on the second task is a direct function of alignment on the first task. We verify this with simulations in Figure 7 (b). Since $K_{y}(T)$ is monotonically increasing with $T$, longer training on the first task will accelerate learning of the reversed task, providing an explanation for overtraining reversal. The key intuition is that adaptation time depends on alignment $K_{y}=K_{-y}$, where $\bm{y}^{\top}\bm{K}\bm{y}=(-\bm{y}^{\top})\bm{K}\bm{(}-y)$, so overtraining leads to faster adaptation.
(a) Pretraining Feature Movement ($t<T$)
(b) Overtraining = Faster Reversal ($t>T$)
Figure 7: A simple two-layer model explains faster reversal learning after overtraining on the original task. (a) We plot the original hidden representations (projected from $N$ hidden dimensions to the top two principal components) of a width $N=250$ network for $4$ odor mixtures, with two odors positive and two odors negatively labeled. (b) When training a new readout on the reversal task $-y(x)$ after pretraining for $t$ steps on the original labels $y(x)$, we find that additional pretraining steps $T$ accelerate learning the reversal task, consistent with theory.
@@ A.1 Methodological and implementation details
The raw data consists of around 70 sessions, each representing a recording of  20 random neurons on a given mouse-day. The data is spikes over time for each neuron on each trial, of which there are 200-300 in a session. After odor onset, mice are given 2.5s to lick in response. We count the number of spikes in the final 500ms part of this window and decode on this number, which gives us a $N\times T$ neurons by trials spike count matrix to decode on for each mouse-day. We drop neurons with low (less than 4) number of spikes across all trials classifying them as unresponsive to the odors, we z-score this matrix and use 10-fold cross validation to get decoding accuracy, over 20 iterations. Since session to session data is very noisy, we plot smoothed statistics using a sliding window and plot standard error of all statistics. It should be noted the last few days of mouse $T$ overtrainining are dropped because of notable inconsistencies in neural data such as the decoding accuracy plummeting to far below chance while behavior remained high, suggesting an unknown and idiosyncratic cause for unreliable decoding accuracy in this mouse.222[Berners-Lee et al., 2023] include these unreliable decoding trials in their analysis. Note that the activity matrix was z-scored across trials, so every neuron contributed equally in our analysis.
@@ A.2.2 Important Experimental Details for General Sensory Discrimination Tasks
A.3 Continued Introduction
A.4 Mathematical Details on Fisher Discriminant, LDA, and Margin
@@ A.5 Representational Separation Ablation
A.6 Hard Margin Ablations
We begin with a review of the Cross-Entropy vs Hinge loss definitions. Let $y\in\{0,1\}$ represent the true class label, and $\hat{y}\in[0,1]$ represent the predicted probability for the positive class (i.e., $P(y=1\mid x)$ ). The cross-entropy loss $L_{\mathrm{CE}}$ for a single example is defined as:
$L_{\mathrm{CE}}(y,\hat{y})=-(y\log(\hat{y})+(1-y)\log(1-\hat{y})).$
In hinge loss, the labels $y$ are typically encoded as $y\in\{-1,+1\}$, and the model outputs a score $f(x)\in\mathbb{R}$, where the sign of $f(x)$ determines the class prediction. For simplicity, let’s assume $f(x)$ is the output of a linear model, $f(x)=w\cdot x+b$.
The hinge loss $L_{\text{hinge }}$ for a single example is defined as:
$L_{\text{hinge }}(y,f(x))=\max(0,C-y\cdot f(x)),$
where $C$ is the hard margin parameter to be specified, which we take as $C=1$. The key point is that in a Hinge objective, loss vanishes after the learned classifier classifies each training example with margin at least $C$, whereas loss does not vanish in this discontinuous way for the Cross-Entropy objective, driving continued late-time learning. We can see in Figure 9that test loss stops improving when train loss does, and this is coincedent with an abrupt ceasation in the increase in margin, in contrst to the continued margin maximization driven by Cross-Entropy in the main text.
Figure 9: Loss function ablation for synthetic model that traces asymptotic margin maximization induced by cross-entropy as the cause for late-time decrease in test loss/late-time separation of representations. “Late-time” refers to when training loss plateaus at a low value.
A.7 Biologically Plausible Model
We now repeat the above experimental setup of our synthetic model with an architecture that takes piriform cortex more seriously [Stern et al., 2018, Krishnamurthy et al., 2022, Mathis et al., 2016].
Our biologically plausible model consists of three main processing layers plus an output layer, with additional feedback connections to model recurrent dynamics. The first layer serves as a sensory input layer, analogous to inputs from the olfactory bulb. It projects the input to a higher-dimensional space, using ReLU activation followed by dropout with probability 0.2 to maintain sparse coding, a key feature of sensory representations in the piriform cortex. The second layer acts as a dense associative layer, modeling the extensive recurrent connectivity between pyramidal cells. This layer maintains the expanded dimensionality from layer one and includes a parallel feedback pathway. The feedback is implemented through a separate linear transformation followed by ReLU activation, and the result is added back to the main pathway via a skip connection. Both the main and feedback pathways use the same dimensionality to preserve the associative nature of this layer. The third layer implements modulatory control, inspired by inhibitory interneuron circuits. This layer reduces dimensionality compared to layer two and uses LeakyReLU activation followed by a higher dropout rate (0.3) to model inhibitory effects. The reduction in dimensionality reflects the modulatory rather than representational role of this layer. The output layer performs a final linear transformation to a single scalar value (since we’re training on binary classification). We train without weight decay and learning rate 1e-2. We also add Gaussian noise to the forward pass to simulate stochastic activity in sensory cortex, hence why there are multiple target odor representations.
Results are shown in Figure 10. The persistence of the phenomenology identified in the main text suggests that the objective function and task structure, rather than architectural details, are the key here.
Figure 10: Architecture ablation for biologically plausible model of piriform cortex trained on the same binary classification task to separate target and nontarget odors. We see loss on held out probe odors continues to decrease after low train loss is achieved, since cross-entropy loss is once again used as the objective.